A kid bought a slushy at Knott's Berry Farm a 9qt. of a 5% sugar solution was mixed with a 11qt. of the 45% sugar solution. What is the concentration of the mixture?
Answer by flame8855(424)

total volume = 9 + 11 =20
0.05*(9) + (11)*(0.45)= 0.45 + 4.95 = 5.4
so 5.4/20 *100 = 27%